OpenStudy (anonymous):

How do I do the integral of sin^(-1)(x)dx?

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin^{-1}x dx\]

OpenStudy (anonymous):

i am going to guess parts

OpenStudy (anonymous):

matter of fact i am almost sure it is parts \[\int u dv =uv-\int vdu\] with \[u=\sin^{-1}(x), du =\frac{1}{\sqrt{1-x^2}}, dv=1dx, v =x\]

OpenStudy (anonymous):

same sort of trick as \[\int \ln(x)dx\]

OpenStudy (anonymous):

you get \[x\sin^{-1}(x)-\int \frac{x}{\sqrt{1-x^2}}dx\] and the second integral is a relatively easy \(u\) sub with \(u=1-x^2, du=-2xdx\) etc

OpenStudy (anonymous):

Oh I see where the parts are now, thank you.